Integration Procedures (Integration by Parts)


∫ x³dx/cuberoot of (8-x²) from 0 to √7


1
Expert's answer
2021-04-13T23:57:29-0400

Integration Procedures (Integration by Parts)


∫ x³dx/cuberoot of (8-x²) from 0 to √7

Solution:

"I=\int_0^{\sqrt7}\frac{x^3}{\sqrt[3]{8-x^2}}dx=\frac12\int_0^{\sqrt7}\frac{x^2}{\sqrt[3]{8-x^2}}dx^2" "=\frac12\int_0^{7}\frac{t}{\sqrt[3]{8-t}}dt"

Integration by parts:

"u=t" ; "du=dt" ;

"dv=\frac{dt}{\sqrt[3]{8-t}}" ; "v=-\frac32(8-t)^{\frac23}" .

"I=-\frac34\cdot t(8-t)^{\frac23}|_0^7+\frac12\int_0^7\frac32(8-t)^{\frac23}dt=-\frac34\cdot7-" "\frac34\cdot\frac35(8-t)^{\frac53}|_0^7=-\frac{21}{4}-\frac{9}{20}+\frac{9\cdot2^5}{20}=\frac{87}{10}=8.7"

Answer: "\int_0^{\sqrt7}\frac{x^3}{\sqrt[3]{8-x^2}}dx=8.7" .